Find the midpoint between these points:
(-3, 7) and (4, -5)

Answer by Alan3354(21584)

(-3, 7) and (4, -5)
(x1 + x2)/2 = (-3+4)/2 = 1/2
----------------
(y1 + y2)/2 = (7 - 5)/2 = 1
Mid point = (1/2,1)